True or False. When non-conservative forces are present, the amount of work done increases with the length of the path.


Use the information below the answer the following 3 questions.

A 50 kg crate is being dragged across a floor by a force of 225 N at an angle of 40o from the horizontal. The crate is dragged a distance of 5.0 m and the frictional force is 60 N.


1) The work done on the crate by the applied force is ___x102 Nm.

2) The work done on the crate by the frictional force is -___x102 Nm.

3) The net work done on the crate is ___x102 Nm.

Do not use rounded answers in subsequent calculations!

1
Expert's answer
2021-07-15T12:41:36-0400

"\text{A-priory:}"

"\text{A non-conservative force is one for which work done on the object}"

"\text{depends on the path taken by the object.}"


"\text{Answer: True}"


"W = Fs \cos\alpha"

"W = 225*5*\cos40\degree=8.62*10^2"

"W_{fr}=-F_{fr}s"

"W_{fr}=-F_{fr}s= 60*5= 3*10^2"

"W_{clear}= (F\cos\alpha -F_{fr})s"

"W_{clear}= (225\cos40\degree -60)*5=5.62*10^2"


"\text{Answer:}"

1) The work done on the crate by the applied force is 8.62x102 Nm.

2) The work done on the crate by the frictional force is -3x102 Nm.

3) The net work done on the crate is 5.62x102 Nm.